OpenStudy (anonymous):

write the series with summation notation. 1/2+1/4+1/6+ . . . +1/40

OpenStudy (tkhunny):

1) 2 = 2*1 2) 4 = 2*2 3) 6 = 2*3 Find a pattern and index it.